OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164 step by step please

OpenStudy (amistre64):

one way is to develop a system of equations ....

OpenStudy (amistre64):

\[a_o+d(19-1)=-58\\ a_o+d(21-1)=-164\]

OpenStudy (amistre64):

might be a1 in in that case ....

OpenStudy (amistre64):

a + 18d = -58 a + 20d = -164 -a - 18d = 58 a + 20d = -164 --------------- (20-18)d = (58-164) d = (58-164)/(20-18)

OpenStudy (anonymous):

so its a + 18d = -58 a + 20d = -164 -a - 18d = 58 a + 20d = -164 --------------- (2)d = (-106) d = (-106)/(2) d=-53?

OpenStudy (anonymous):

@amistre64 hey i'm still lost as in what to do next

OpenStudy (amistre64):

you found the value for d, use it to find a; then all the parts will be accounted for

OpenStudy (amistre64):

a + 18d = -58 a =-58- 18d

OpenStudy (amistre64):

im assuming you know the general form of an arithmetic sequence to start with

OpenStudy (anonymous):

i really don't , this is a new lesson for me and i'm struggling to understand. a=-58-18(-53) a=-58-(-954) a=-58+954=896

OpenStudy (amistre64):

the general form of defining the nth term of an arithmetic sequence is:\[a_1+d(n-1)\] hence the need to define the a and d parts

OpenStudy (amistre64):

we are given two solutions; which we can then create 2 equations with 2 unknowns with; which is what we started out doing; then we determine their values

OpenStudy (amistre64):

test it out; does: 896 - 53(19-1) = -58 ? does: 896 - 53(20-1) = -164 ?

OpenStudy (amistre64):

if so; then:\[a_n=896-53(n-1)\]

OpenStudy (amistre64):

21-1 that is; fingers got a little ahead

OpenStudy (anonymous):

896-53(18)=896-954=-58 896-53(20)=896-1060=-164 ok it makes sense now thank you

OpenStudy (amistre64):

youre welcome